A gold doubloon 6.1 cm in diameter and 2.0 mm thick is dropped over the side of a pirate sheep. When it comes to rest on the ocean floor at a depth of 770 m, how much has its volume changed?​


1
Expert's answer
2021-11-29T11:42:15-0500

Assume the water pressure compresses the coin only in vertical direction along the axis of the cylinder.

The pressure:


"p=\rho gh."

The force that compresses the coin:


"F=pA."

The stress in the material of the coin caused by the applied force:


"\sigma=\frac FA=p."

The strain:


"\epsilon=\frac{\Delta h}h=\frac p E,\\\space\\
\Delta h=h\frac pE=\frac{\rho gh^2}{E}."

The volume change is


"\Delta V=\Delta hA,\\\space\\
\Delta V=\frac{\rho \pi gh^2d^2}{4E}=1.5·10^{-15}\text{ m}^3."

Note: we took sea water density as 1030 kg/m3, the young's modulus of gold is 79 GPa.